OpenStudy (anonymous):

Can anyone help with an assignment on Basic Rules of Differentiation / Marginal Analysis?

OpenStudy (anonymous):

2) \[g(x)=y=x^x\] \[\ln y=\ln x^x\]

OpenStudy (anonymous):

\[\ln y=x \ln x\] \[\frac1y y'=1 \ln x+x \frac1x\] \[y'=y(\ln x+1)\] \[y'=x^x(\ln x+1)\]

OpenStudy (anonymous):

Is there another way of figuring this out without using ln? We haven't done that in any assignments yet.

OpenStudy (anonymous):

\[y'=g'(x)=x^x (\ln x+1)\] \[g'(2)=2^2(\ln2+1)\] \[g'(2)\approx6.77\]

OpenStudy (anonymous):

3) f(x)=ax^2+bx+c f(6)=150 f(10)=50

OpenStudy (anonymous):

f(8)=100

OpenStudy (anonymous):

150=36a+6b+c 50=100a+10b+c 100=64a+8b+c